OpenStudy (anonymous):

y varies directly with x. When y = 12, x = 16. Find x when y = 24.

OpenStudy (australopithecus):

y = 12 when x = 16 thus we have the point (16,12) and we have the point (x, 24) x/16 = 24/12 x = 24(16)/12 x = 32

OpenStudy (australopithecus):

I just used a ratio x2/x1 = y2/y1

OpenStudy (australopithecus):

since they change proportionally the two ratios are equal